THEORETICAL INVESTIGATION OF SPIN-ORBIT VIBRONIC COUPLING EFFECTS IN THE ELEC-TRONIC GROUND STATE OF CrCN

\underline{ILIAS SIOUTIS} and WOLFGANG DOMCKE, Department of Chemistry, Technical University of Munich, D-85747 Garching, Germany.

The spin-orbit-induced vibronic coupling interactions in the ^{6}Σ^{+} electronic ground state of the linear CrCN molecule are investigated, employing the microscopic (Breit-Pauli) spin-orbit (SO) coupling operator. The 6×6 Hamiltonian matrix is derived in a diabatic spin-orbital electronic basis set including terms up to second order in the expansion of the molecular Hamiltonian in the bending normal coordinate. Ab initio calculations of the potential energies of the ^{6}Σ^{+} state are performed as a function of the bending normal coordinate. The fitting of the spin degeneracy of the ^{6}Σ^{+} state via various SO coupling terms is investigated. The predicted electronic structure of the X̃ ^{6}Σ^{+} electronic state of CrCN shows a good agreement with the experimentally^{a} determined Cr-C bond length but less so for the C-N bond length. The nature of the metal-ligand bonding is also discussed.

aM. A. Flory, R. W. Field and L. M. Ziurys Mol. Phys. \underline{105}, 585 (2007).